OpenStudy (anonymous):

which method would be the simplest way to solve the system? A. graphing b.substitution c. elimination d. distributive

OpenStudy (anonymous):

please help!! @allie_bear22 @tvartheman @SeaTurtle113

OpenStudy (seaturtle113):

It depends on what the system is, but I think the answer is elimination.

OpenStudy (anonymous):

sorry y=1/2x and 2x+3y=28 @SeaTurtle113

OpenStudy (anonymous):

I need at least an explaination plz

OpenStudy (haseeb96):

substitution put y in the 2x+3y-28 then you get 2x+3(1\2x)=28 4x+3\2x =28 4x+3=56x 3=56x-4x 3=52x x=52\3 put x in the y=1\2x y=1\2(52\3) y=3\104

OpenStudy (anonymous):

thank you:)

OpenStudy (campbell_st):

the method is correct... but the solution is incorrect quick question is the 1st equation \[1. .. y = \frac{1}{2} x\] or 2. \[y = \frac{1}{2x}\]

OpenStudy (campbell_st):

lol... have a look at @Haseeb96 solution he has \[2x + 3(1/2x) = 28\] when then becomes \[4x + 3/2x = 28\] where did 4 x come from then there is the major mistake multiplying every 2x, using his equation should be \[4x \times 2x + 3 = 56x... or.... 8x^2 + 3 = 28x\] and not \[4x + 3 = 56x\]

OpenStudy (anonymous):

Thaanks:) @cambell_st